OpenStudy (anonymous):

what is f(x + 1) when f(x) = |x| + 4 ? i got x - 5 but i am not sure

OpenStudy (anonymous):

you get \[f(x+1)=|x+1|+4\]

OpenStudy (amistre64):

x-5 is not it

OpenStudy (anonymous):

that is replace x by x + 1

OpenStudy (amistre64):

you mighta typoed it tho, x+5 is good

OpenStudy (anonymous):

\[f(\spadesuit)=|\spadesuit|+4\]

OpenStudy (anonymous):

god but i love that!

OpenStudy (anonymous):

can you explain why it is x + 5? because if you subtract 1 from the absolute value you should get x + 3...

OpenStudy (anonymous):

\[f(\heartsuit + \diamondsuit)=|\heartsuit + \diamondsuit|+1\]

OpenStudy (anonymous):

it isn't!

OpenStudy (anonymous):

whats with the weird characters...

OpenStudy (anonymous):

replace x by x + 1. that is it

OpenStudy (anonymous):

you cannot simplify more?

OpenStudy (anonymous):

you get \[|x+1|+4\] which is not \[x+5\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i= |x%2B1|%2B4 ; x+5 is good as long as x is positive

OpenStudy (anonymous):

|x| + 5, is the correct answer

OpenStudy (amistre64):

that aint right

OpenStudy (anonymous):

f(x+1) = |x+1| + 4 = |x| +|1|+4 = |x| + 5

OpenStudy (amistre64):

-(x+1)+4 and (x+1)+4 maybe

OpenStudy (amistre64):

i dont think you can seperate a unary function like that

OpenStudy (anonymous):

yeah, ur right

OpenStudy (amistre64):

what does the title of the window say when you in here?

OpenStudy (amistre64):